Let the price elasticity of demand be 1. Let the price be $100/MWh. Let consumption be 50,000MWh. Which of the following would be the demand equation for electricity?

Q = 500P + 1000

Q = 500P

Q = 5000P - 500,000

None of the above.
Answer #1

Putting P=100, we getQ=50,100 in first option, for second we get Q=500*100=50,000 for third option we get Q=5000*100-500,000=0 Since actual consumption is 50,000 only option two fits and dQ/dP for option 2 is also 1 which is elasticity Hence, Q=5000 P is correct